# DecoderTracker: Decoder-Only Method for Multiple-Object Tracking

### v4

@@ 4.3 Performance Evaluation and Comparison
In the MOT17 (Table 3), DecoderTracker demonstrates competitive performance, particularly in handling complex tracking scenarios, with competitive scores in HOTA and AssA. Despite trailing slightly behind MOTR in these metrics, DecoderTracker compensates with a comparable DetA and excels in IDF1 by almost 10 points. This advantage can be attributed to DecoderTracker’s simplified network structure, which enables competitive tracking accuracy with reduced training data demand. Additionally, DecoderTracker maintains robust overall performance, as evidenced by its high MOTA score, highlighting its suitability for real-world tracking applications.
@@
4.4 Ablation Study
To assess the proposed training strategy and the efficacy of TBSP, ablation experiments were conducted on the Dancetrack benchmark. The training set includes the entire Dancetrack training data, with testing involving 25 sequences from the validation set. Additionally, the impact of two methods for generating query positions—utilizing the sinusoidal functions for encodings from DAB-DETR[48] or employing an MLP, on the model’s performance was investigated.
@@ 4.4 Ablation Study
Training Stage HOTA AssA MOTA
First 38.5 21.6 76.8
Second 44.5 25.5 84.2
Final (TBSP) 51.9 34.6 84.9
Final 53.6 35.9 84.7
(b) The presence or absence of TAN during the second training stages can significantly impact the performance of the model.
Model HOTA AssA MOTA
Second 44.5 25.5 84.2
Second(w/o) 42.1 21.4 80.8
Final 53.6 35.9 84.7
Final(w/o) 53.4 36.1 84.5
(c) The impact of $\delta_{iou}$ on model performance when using TBSP
$\delta_{iou}$ HOTA AssA MOTA
0.3 49.7 32.3 79.1
0.5 51.9 34.6 84.9
0.7 51.3 34.8 83.2
0.9 50.0 32.8 81.3
@@
The performance of DecoderTracker improves consistently across different training stages, peaking in the final stage. The choice of $\delta_{iou}$ impacts precision and recall trade-offs, with higher values leading to a slight decrease in performance. Position generation strategies, MLP and sinusoidal functions, yield similar performance in terms of detection accuracy, with slight differences in FPS. Overall, the ablation studies demonstrate the effectiveness of DecoderTracker across different training conditions and model configurations, showcasing its adaptability and versatility in MOT field.
@@
It is noteworthy that in Table 5(b), we demonstrate the impact of the presence or absence of TAN during the second phase of training on the network’s tracking performance at different stages. It can be observed that the presence or absence of TAN in the second phase has almost no effect on the final network performance. However, the absence of TAN does impact the tracking performance during the second phase, which also validates the effectiveness of our proposed TBSP strategy in enhancing tracking capabilities.
@@ 5 Conclusion and Limitation
This paper presents an efficient decoder-only end-to-end MOT (Multi-Object Tracking) model that boasts faster inference speeds and reduced training times. Additionally, this paper introduces new training strategies aimed at accelerating model convergence. Although device limitations prevented us from validating our training strategies on the original MOTR model, we believe these strategies are effective. This research provides a more efficient benchmark model for the multi-object tracking field compared to MOTR, laying a foundation for the emergence of real-time end-to-end MOT models and offering insights for future development. In subsequent work, we should further explore more efficient network architectures to construct real-time MOT models. Of course, we also plan to incorporate techniques from MOTRv2, MOTRv3, and CO-MOT into DecoderTracker to enhance tracking performance.
Appendix A More Discusstion About TBSP
Figure 5: Refinement process for DecoderTracker on partially trained track queries. When DecoderTracker undergoes comprehensive training with the TALA strategy, TBSP will no longer be utilized. At that point, the track query generation method of DecoderTracker aligns with MOTR, as illustrated in the Fig. 1.
Setting HOTA MOTA Epochs Total Time
TBSP 53.6 84.7 27 56.7 hours
w/o 53.4 84.2 40 85.2 hours
The initial reason for developing the MO-YOLO model was not due to slow inference speed, but rather the extensive training resources required. Using eight V100 GPUs with pretrained weights, training on an extended MOT17 dataset, which contained approximately 10,000 images, required about 84 hours for 200 epochs. Each epoch took around 0.42 hours, with each GPU processing about 3,000 images per hour.
Fortunately, our DecoderTracker model has addressed this issue. Although its inference speed is not yet real-time, its convergence rate is already satisfactory. Although we have not achieved real-time inference speeds, we have significantly improved the speed of MOTR.
We believe the effectiveness of the TBSP strategy may be due to its function as a form of weak or self-supervised training for motion features, compared to TALA. Many examples demonstrate that undergoing weakly supervised training before strongly supervised training can enhance performance and speed up convergence. For instance, in the NLP field, large language models like BERT[49] and GPT[1] initially train on a vast amount of unlabeled or weakly labeled data to learn basic language patterns. Subsequently, these models undergo strong supervision fine-tuning for specific tasks like sentiment analysis or question answering. This approach has been empirically proven to significantly boost the performance of LLMs.
In our field of computer vision, this technique is also quite common, with the most classic example being the SimCLR series[50, 51]. Typically, these models are pretrained on unlabeled data to learn image features, then fine-tuned on strongly supervised tasks such as classification or detection.
We believe that the TBSP, a weak supervision process, enables the model to preliminarily learn object motion information that would otherwise require many more training iterations. It’s akin to learning to ride a bicycle where training wheels help beginners to ride without obstacles while also giving them a feel for biking, which aids in eventually mastering the skill. A similar situation may exist here.
Although TBSP did not enhance model performance, it reduced the number of iterations needed for convergence, which might be considered its greatest contribution.
Appendix B Design of IMFENet
The design of this feature network is intended to enhance the model’s efficiency and increase its inference speed. Thus, our design primarily references the network structure of YOLOv8. The backbone structure of our model is almost identical to that of YOLOv8. Our main modification involves changes to the neck structure in YOLOv8, to ensure that its output aligns with the input requirements of the decoder.
Figure 6: The difference between our proposed model and YOLOv8 in terms of modules. Our modules are more lightweight. Specific model details can refer to the figure.3.
@@ Appendix B Design of IMFENet
Table 7: The impact of our network and YOLOv8 structure on DecoderTracker performance on the Dancetrack val set. Training time does not include the first stage
Setting HOTA IDF1 MOTA AssA FPS
IMFENet 53.6 53.8 84.7 35.9 19.6
YOLOv8 52.4 51.4 84.2 36.1 18.9
Appendix C Limitation
Undoubtedly, the greatest limitation of DecoderTracker, similar to that of MOTR, is its performance on the MOT17 benchmark, which does not match that of tracking-by-detection models. Both models fail under similar circumstances, as illustrated in Figure.7. This issue may, like with MOTR, stem from the insufficient data volume on the MOT17 benchmark, which does not adequately train the model. Clearly, despite this limitation, DecoderTracker still represents a significant improvement over MOTR.

### v5

@@ 4.3 Main Result
In the MOT17 benchmark (Table 3), DecoderTracker shows strong performance, especially in complex tracking scenarios. While slightly behind MOTR in HOTA and AssA, it compensates with comparable DetA and excels in IDF1 by nearly 10 points. This advantage is due to its simplified network structure, which achieves competitive accuracy with less training data. Moreover, DecoderTracker maintains robust overall performance, as shown by its high MOTA score, making it suitable for real-world applications. When additional training data is introduced, DecoderTracker’s performance further improves. For instance, DecoderTracker* achieves a significant boost in IDF1 to 77.2, HOTA to 61.8, and MOTA to 76.8. Similarly, DecoderTracker+* also shows substantial gains. This clearly demonstrates the positive impact of more training data on DecoderTracker’s tracking capabilities, highlighting its potential for even better performance with increased data support.
@@
Figure 8: The visualisation results of our proposed DecoderTracker on various datasets show that DecoderTracker performs well both under long time series and short time series, and also has good tracking performance for different kinds of objects. The visualization results of DecoderTracker and DecoderTracker+ are essentially identical, so we won’t display the results of D+ here.
4.4 Ablation Study Of DecoderTracker
To assess the proposed training strategy and the efficacy of TBSP, ablation experiments were conducted on the Dancetrack benchmark. The training set includes the entire Dancetrack training data, with testing involving 25 sequences from the validation set. Additionally, the impact of two methods for generating query positions—utilizing the sinusoidal functions for encodings from DAB-DETR [31] or employing an MLP, on the method’s performance was investigated.
@@ 4.4 Ablation Study Of DecoderTracker
Training Stage IDF1 HOTA MOTA
First 38.6 38.5 76.8
Second 44.3 44.5 84.2
Final (TBSP) 51.7 51.9 84.9
Final 53.4 53.6 84.7
(b) The impact of $\delta_{iou}$ on method performance when using TBSP
$\delta_{iou}$ IDF1 HOTA MOTA
0.3 50.1 49.7 79.1
0.5 51.7 51.9 84.9
0.7 51.3 51.3 83.2
0.9 49.8 50.0 81.3
(c) The presence or absence of TAN during the second training stages can significantly impact the performance of the method.
method IDF1 HOTA MOTA
Second 44.4 44.5 84.2
Second(w/o) 41.9 42.1 80.8
Final 53.4 53.6 84.7
Final(w/o) 53.2 53.4 84.5
@@
The performance of DecoderTracker improves consistently across different training stages, peaking in the final stage. The choice of $\delta_{iou}$ impacts precision and recall trade-offs, with higher values leading to a slight decrease in performance. Position generation strategies, MLP and sinusoidal functions, yield similar performance in terms of detection accuracy, with slight differences in FPS. Overall, the ablation studies demonstrate the effectiveness of DecoderTracker across different training conditions and method configurations, showcasing its adaptability and versatility in MOT field.
@@
It is noteworthy that in Table 5(c), we demonstrate the impact of the presence or absence of TAN during the second phase of training on the network’s tracking performance at different stages. It can be observed that the presence or absence of TAN in the second phase has almost no effect on the final network performance. However, the absence of TAN does impact the tracking performance during the second phase, which also validates the effectiveness of our proposed TBSP strategy in enhancing tracking capabilities.
4.4.1 More Discusstion About TBSP
The initial reason for developing the DecoderTrack method was not due to slow inference speed, but rather the extensive training resources required. Using eight V100 GPUs with pretrained weights, training on an extended MOT17 dataset, which contained approximately 10,000 images, required about 84 hours for 200 epochs. Each epoch took around 0.42 hours, with each GPU processing about 3,000 images per hour.
Fortunately, our DecoderTracker method has addressed this issue. Although its inference speed is not yet real-time, its convergence rate is already satisfactory. Although we have not achieved real-time inference speeds, we have significantly improved the speed of MOTR.
We believe the effectiveness of the TBSP strategy may be due to its function as a form of weak or self-supervised training for motion features, compared to TALA. Many examples demonstrate that undergoing weakly supervised training before strongly supervised training can enhance performance and speed up convergence. For instance, in the NLP field, large language methods like BERT[53] and GPT[54] initially train on a vast amount of unlabeled or weakly labeled data to learn basic language patterns. Subsequently, these methods undergo strong supervision fine-tuning for specific tasks like sentiment analysis or question answering. This approach has been empirically proven to significantly boost the performance of LLMs.
Setting IDF1 HOTA Epochs Total Training Time
TBSP 54.2 53.6 27 56.7 hours
w/o 53.5 53.4 40 85.2 hours
In our field of computer vision, this technique is also quite common, with the most classic example being the SimCLR series[12, 13]. Typically, these methods are pretrained on unlabeled data to learn image features, then fine-tuned on strongly supervised tasks such as classification or detection.
We believe that the TBSP, a weak supervision process, enables the method to preliminarily learn object motion information that would otherwise require many more training iterations. It’s akin to learning to ride a bicycle where training wheels help beginners to ride without obstacles while also giving them a feel for biking, which aids in eventually mastering the skill. A similar situation may exist here.
Although TBSP did not enhance method performance, it reduced the number of iterations needed for convergence, which might be considered its greatest contribution.
4.5 Ablation Study Of DecoderTracker+
As can be seen from the results in Tables 2, 3, and 4, the performance of DecoderTracker and DecoderTracker+ is almost the same. This is not difficult to understand. In Section 3.3, we discussed that empty queries do not affect the network’s output results after reasonable mask generation. DecoderTracker+ does not improve the network; instead, it addresses the optimization issues caused by dynamic data. At the same time, to better illustrate the effectiveness of the generated masks, we provided the performance of DecoderTracker+ with and without masks on the Dancetrack val set in Table 7. The results show that when masks are not used, the inactive queries still have an impact on the final tracking results. Combining this with previous data, both theoretical and practical evidence demonstrates that the use of masks does not affect tracking performance when FSQM is increased after reasonable mask generation.
Table 7: The impact of mask on the performance of DecoderTracker+ on the Dancetrack val set.
mask IDF1 HOTA MOTA AssA
53.6 53.4 84.9 36.0
w/o 46.2 45.8 78.3 31.7
4.5.1 Experimental Analysis of FENet
Figure 9: The difference between our proposed method and YOLOv8 in terms of modules. Our modules are more lightweight. Specific method details can refer to the Figure.3. Additionally, it should be noted that there is not much difference between this part of YOLO11 and YOLOv8, so no comparison has been made
Table 8: The impact of our network and YOLOv8 structure on DecoderTracker and DecoderTracker+ performance on the Dancetrack val set.
Setting IDF1 HOTA MOTA AssA FPS
DecoderTracker:
FENet 53.8 53.6 84.7 35.9 19.6
YOLOv8 51.4 52.4 84.2 36.1 18.9
1.88 DecoderTracker+:
FENet 53.6 53.4 84.9 35.7 28.2
YOLOv8 52.1 53.2 85.4 36.8 26.7
@@ 5 Conclusion and Limitation
This paper proposes an efficient decoder-only end-to-end MOT method, which achieves faster inference speed and shorter training time compared to the baseline MOTR. Additionally, we introduce new training strategies aimed at accelerating the convergence of the method. Due to hardware limitations, we were unable to verify our training strategies on MOTR, but we believe these strategies are generalizable. Finally, we conducted a detailed and in-depth analysis of the slow inference process in MOTR, finding that dynamic data significantly impacts inference speed. Subsequently, we applied FSQM to address this issue, greatly enhancing the inference speed of the tracking-by-query end-to-end paradigm. Overall, this study provides a more efficient benchmark method for the multi-object tracking field, lays the foundation for the emergence of real-time end-to-end MOT methods, and offers insights for future development. In future work, we should further explore more efficient network architectures to build real-time MOT methods. Of course, we also plan to integrate the technologies from MOTRv2, MOTRv3, and CO-MOT into DecoderTracker to improve tracking performance.
Undoubtedly, the greatest limitation of DecoderTracker, similar to that of MOTR, is its performance on the MOT17 benchmark, which does not match that of tracking-by-detection methods. Both methods fail under similar circumstances, as illustrated in Figure.10. This issue may, like with MOTR, stem from the insufficient data volume on the MOT17 benchmark, which does not adequately train the method. Clearly, despite this limitation, DecoderTracker still represents a significant improvement over MOTR. Although we attempted to use a mixed dataset to address this issue, our method still underperforms compared to SOTA tracking-by-detection methods. How to resolve this issue will also be a key focus of future research.